what is the area of a triangle if one leg is 2 root 3 ft. and the hypotenuse is 4 ft.
Answer by stanbon(75887)

what is the area of a triangle if one leg is 2 root 3 ft.
and the hypotenuse is 4 ft.
--------
Draw the triangle ; label the base 2sqrt(3) ft ;
label the hypotenuse 4 ft.
--------
Solve for the height::
height = sqrt(4^2-(2sqrt(3))^2)
-----
height = sqrt(16-12)
height = 2 ft
-----
Area = (1/2)base*height = (1/2)(2sqrt(3))(2) = 2sqrt(3) sq.ft.
==================
Cheers,
Stan H.
==================